OpenStudy (anonymous):

What is the value of x? 5.1 in 4.9 in 4 in 4.5 in i think its 4.9 but not sure

OpenStudy (hba):

Where is the diagram ?

OpenStudy (hba):

Use similar triangles. Ratio for sides. \[4/(4+18) = x / (x+22)\]

OpenStudy (anonymous):

5.5=22x/x?

OpenStudy (hba):

What ?

OpenStudy (anonymous):

i'm so confused

OpenStudy (hba):

Cross multiply \[4(x+22)=x(4+18)\]

OpenStudy (hba):

Now can you solve for x ?

OpenStudy (anonymous):

so its 5.1

OpenStudy (hba):

Did you calculate it ?

OpenStudy (anonymous):

yeah... i got 5.1

OpenStudy (hba):

Show me your working...

OpenStudy (hba):

@xxjakkixx

OpenStudy (anonymous):

wait i just redid it and i got 4.... i dont understand this at all

OpenStudy (hba):

I said show me your work;

OpenStudy (hba):

You have to use the distributive property to solve it a(x+y)=ax+ay